Ron tossed unbiased coin, on each toss, he wins P50 if head appears, and loses P40 if tail appears. What will be the expected value of his gain? 


1
Expert's answer
2022-03-02T10:04:32-0500

Solution:

Expected value"=50\times \dfrac12-40\times \dfrac12=25-20=5"

Expected value of gain is P5.